OpenStudy (anonymous):

In probabilityland, one third of population is senior. Every winter 20% of population gets the flu. on average, two out of five seniors get the flu. what is the probability that a person with the flu is senior?

OpenStudy (anonymous):

here is what I did

OpenStudy (anonymous):



OpenStudy (anonymous):

"that's probability of senior with flu" whereas they are asking for "probability of person with a flu being a senior" I am not sure if they are same

OpenStudy (anonymous):

@zarkon

OpenStudy (anonymous):

or let event A be senior event B be people with flu A=1/3 S B=.2 S 2/5B=(A ∩ B) probability of people with flu being a senior (A ∩ B) ------------ B

OpenStudy (zarkon):

so what is your answer

OpenStudy (anonymous):

well, I thought it was 2/15 but I am having second thoughts

OpenStudy (zarkon):

you want P(senior|flu) correct?

OpenStudy (anonymous):

yes, probability of people with flu who are senior

OpenStudy (zarkon):

\[P(S|F)=\frac{P(F|S)P(S)}{P(F)}\]

OpenStudy (anonymous):

so , P(F|S) is what I found to be 2/15?

OpenStudy (zarkon):

\[P(F|S)=\frac{2}{5}\]

OpenStudy (anonymous):

ok, also is this a defination \[P(S|F)=\frac{P(F|S)P(S)}{P(F)}\]

OpenStudy (zarkon):

\[P(A|B)=\frac{P(A\cap B)}{P(B)}\] is the definition of conditional probability

OpenStudy (zarkon):

use this definition twice to get the formula I have above

OpenStudy (zarkon):

\[P(A|B)=\frac{P(A\cap B)}{P(B)}\] \[P(B|A)=\frac{P(A\cap B)}{P(A)}\Rightarrow P(B|A)P(A)=P(A\cap B) \] then \[P(A|B)=\frac{P(A\cap B)}{P(B)}=\frac{P(B|A)P(A)}{P(B)}\]

OpenStudy (anonymous):

wow, thank you very much

OpenStudy (zarkon):

np

OpenStudy (zarkon):

do you have a new final answer?

OpenStudy (anonymous):

(2/5)(1/3) --------- = (1/5) 2/15 ----- = 1/5 2/15 * 5/1=2/3